# The asymptotic Plancherel formula and Lusztig’s asymptotic algebra for $\tilde{\mathsf{A}}_{n}$

### v1

@@ 7.3 The ring $\mathfrak{C}_{\lambda}$ and Lusztig’s asymptotic algebra
Fix bijections $\phi_{ij}:\Gamma_{i}\cap\Gamma_{j}^{-1}\to\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$, as in [31, §2.3]. Write $E_{ij}(a)$ for the matrix with $a$ in the $(i,j)$-position and 00 elsewhere. The following result of Xi reduces the understanding of $\mathcal{J}_{\lambda}$ to understanding the ring $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$.
Theorem 7.7 ([31, Theorem 2.3.2]).
@@ 2. $(b)$
By Theorem 7.2 we have $\mathfrak{c}_{\lambda}(d_{i})=E_{ii}$. Let $w_{ij}=\phi^{-1}_{ij}(\mathsf{w}_{\lambda^{\prime}})\in\Gamma_{i}\cap\Gamma_{j}^{-1}$. By [31, Lemma 2.3.1] we have $w_{ii}=d_{i}$ and $w_{ji}=w_{ij}^{-1}$. Since $E_{ij}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})E_{ji}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})=E_{ii}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}}^{2})=E_{ii}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})$ it follows (under the isomorphism in Theorem 7.7(2)) that $\mathsf{t}_{w_{ij}}\mathsf{t}_{w_{ij}^{-1}}=\mathsf{t}_{d_{i}}.$ By Theorem 7.2 we have $\mathfrak{c}_{\lambda}(w_{ij})=a(\zeta_{\lambda})E_{ij}$ for some $a(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]$. Then the above equation, and the isomorphism from Theorem 5.8, gives (using (5.2)) $E_{ii}=\mathfrak{c}_{\lambda}(d_{i})=\mathfrak{c}_{\lambda}(w_{ij})\mathfrak{c}_{\lambda}(w_{ij}^{-1})=a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})E_{ii},$ and hence $a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})=1$. Writing $a(\zeta_{\lambda})=\sum_{\gamma}c_{\gamma}\zeta_{\lambda}^{\gamma}$ (with $c_{\gamma}\in\mathbb{Z}$ and the sum over $\gamma\in P^{(\lambda)}$) it follows that $\sum_{\gamma}c_{\gamma}^{2}=1$, and hence $a(\zeta_{\lambda})=\epsilon_{ij}\zeta_{\lambda}^{\gamma_{ij}}$ for some $\gamma_{ij}\in P^{(\lambda)}$ and $\epsilon_{ij}=\pm 1$. Thus the elements $w_{ij}$ satisfy $\mathfrak{c}_{\lambda}(w_{ij})=\epsilon_{ij}\zeta_{\lambda}^{\gamma_{ij}}E_{ij},$ with $\epsilon_{ji}=\epsilon_{ij}$ and $\gamma_{ji}=-\gamma_{ij}$ (with $\epsilon_{ii}=1$).
@@
and so $\mathfrak{c}_{\lambda}(\mathsf{w}_{1i})A\mathfrak{c}_{\lambda}(\mathsf{w}_{j1})=a(\zeta_{\lambda})E_{11}$, and hence $a(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$. So $\mathcal{R}_{ij}\subseteq\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$. On the other hand,
$D^{-1}\mathfrak{c}_{\lambda}(w_{i1})\mathfrak{c}_{\lambda}(\mathsf{m}_{\gamma})\mathfrak{c}_{\lambda}(w_{1j})D=\mathfrak{s}_{\gamma}(\zeta_{\lambda})E_{ij},$
and so $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}\subseteq\mathcal{R}_{ij}$, completing the proof of the claim. Thus $D^{-1}\mathfrak{C}_{\lambda}D=\mathrm{Mat}_{N_{\lambda}}(\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}})$.
@@
and so $\mathfrak{c}_{\lambda}(\mathsf{w}_{1i})\mathfrak{c}_{\lambda}(w)\mathfrak{c}_{\lambda}(\mathsf{w}_{j1})=\epsilon_{w}\mathfrak{c}_{\lambda}(\mathsf{m}_{h_{\lambda}(w)})$. Thus $\mathsf{t}_{\mathsf{w}_{1i}}\mathsf{t}_{w}\mathsf{t}_{\mathsf{w}_{j1}}=\epsilon_{w}\mathsf{t}_{\mathsf{m}_{h_{\lambda}(w)}}.$ It is known that the structure constants of $\mathcal{J}_{\lambda}$ are nonnegative (see [16, (3.1.1)]), and hence $\epsilon_{w}=1$. Thus we have $D^{-1}\mathfrak{c}_{\lambda}(w)D=\mathfrak{s}_{h_{\lambda}(w)}(\zeta_{\lambda})$ for all $w\in\Delta_{\lambda}$. ∎
@@
Theorem 7.8 gives rise to a balanced system of cell representations, in the sense of [12, Definition 1.5].
@@
We verify properties $\mathbf{B1}$$\mathbf{B6}$ from [12, Definition 1.5]. Property $\mathbf{B1}$ is the killing property (Theorem 4.3), and property $\mathbf{B2}$ is boundedness (Theorem 4.7, with bounds $\mathbf{a}_{\lambda}=\ell(\mathsf{w}_{\lambda^{\prime}})$). Property $\mathbf{B6}$ (monotonicity of the bounds) is then immediate from Lemma 1.3. Property $\mathbf{B3}$ is the statement that $\pi_{\lambda}$ recognises precisely $\Delta_{\lambda}$ (Theorem 5.7), and property $\mathbf{B4}$ (freeness of the leading matrices over $\mathbb{Z}$) follows from the freeness of the Schur functions and Theorem 7.8. Finally, if $z\in\Delta_{\lambda}$ then $z\in\Gamma_{i}\cap\Gamma_{j}^{-1}$ for some $1\leq i,j\leq N_{\lambda}$ and by Theorem 7.8 we have $\mathfrak{c}_{\lambda}(z)\mathfrak{c}_{\lambda}(d_{j})=\mathfrak{c}_{\lambda}(z)$, verifying $\mathbf{B5}$. ∎
@@
Conjugation by $D$ in the above theorem amounts to choosing a basis associated to a (signed) “fundamental domain” for the action of $\mathsf{T}_{\lambda}$ on $\mathcal{A}_{\lambda}$ different from the standard fundamental domain ${{}^{\lambda}}W$. See [10, Section 4.5] for details. Specifically, the basis chosen is $\{\xi_{\lambda}\otimes\epsilon_{i}X_{\tau_{\gamma_{i}}^{-1}u_{i}}\mid 1\leq i\leq N_{\lambda}\}$ where ${{}^{\lambda}}W=\{u_{i}\mid 1\leq i\leq N_{\lambda}\}$ is ordered as in Theorem 7.2.
@@
$w_{12}$ $=s_{1}s_{3}s_{2}$ $w_{13}$ $=s_{1}s_{3}s_{0}\sigma$ $w_{14}$ $=s_{1}s_{3}s_{2}\sigma$
$w_{15}$ $=s_{1}s_{3}\sigma$ $w_{16}$ $=s_{1}s_{3}s_{2}\sigma^{2}$ $w_{23}$ $=s_{2}s_{1}s_{3}s_{0}\sigma$
$w_{24}$ $=s_{2}s_{1}s_{3}s_{2}\sigma$ $w_{25}$ $=s_{2}s_{1}s_{3}\sigma$ $w_{26}$ $=s_{2}s_{1}s_{3}s_{2}\sigma^{2}$
$w_{34}$ $=s_{3}s_{2}s_{0}s_{1}$ $w_{35}$ $=s_{3}s_{2}s_{0}$ $w_{36}$ $=s_{3}s_{2}s_{0}s_{1}\sigma$
$w_{45}$ $=s_{1}s_{2}s_{0}$ $w_{46}$ $=s_{1}s_{2}s_{0}s_{1}\sigma$ $w_{56}$ $=s_{2}s_{0}s_{1}\sigma$
@@ 2. $(2)$
It is sufficient to prove (1), because (2) follows by applying (1) to $(t_{\gamma}u)^{-1}=t_{-u^{-1}\gamma}u^{-1}$. By counting hyperplanes separating $\mathcal{A}_{0}$ from $t_{\gamma}u\mathcal{A}_{0}$, we have (see [22, (2.4.1)])
@@ A.4 Monotonicity of $\ell(\mathsf{m}_{\gamma})$ with respect to $\preccurlyeq_{\lambda}$
and similarly for $L(\gamma^{\prime})$. Write $\gamma^{\prime}+Q_{\lambda}=\gamma+\epsilon+Q_{\lambda}$ with $\epsilon\in Q^{\lambda}_{+}$. From Lemma 2.16 we have $\epsilon=\sum_{i=1}^{r(\lambda)}d_{i}\tilde{e}_{i}$ where $\sum_{i\in B_{p}}d_{i}=0$ for all $1\leq p\leq r$ and $d_{1}+\cdots+d_{i}\geq 0$ for all $1\leq i\leq r(\lambda)$. Thus $\epsilon=\epsilon_{1}+\cdots+\epsilon_{r}$ where $\epsilon_{p}=\sum_{i\in B_{p}}d_{i}\tilde{e}_{i}\in Q^{\lambda}_{+}$, and for any $1\leq p\leq r$ we have $\gamma+\epsilon_{p}+Q_{\lambda}\in(P/Q_{\lambda})_{+}$ with $\gamma+Q_{\lambda}\preccurlyeq_{\lambda}\gamma+\epsilon_{p}+Q_{\lambda}\preccurlyeq_{\lambda}\gamma^{\prime}+Q_{\lambda}$. Thus it suffices to prove the result in the case $\gamma^{\prime}+Q_{\lambda}=\gamma+\epsilon_{p}+Q_{\lambda}$ for some $1\leq p\leq r$ (that is, $\gamma+Q_{\lambda}$ and $\gamma^{\prime}+Q_{\lambda}$ only differ in one block), and then by well known properties of the dominance order on the symmetric group (see, for example, [29, Corollary 2.7]) it suffices to prove the result when $\epsilon_{p}=\tilde{e}_{i}-\tilde{e}_{j}$ for some $i,j\in B_{p}$ with $i<j$.

### v2

@@ 7.3 The ring $\mathfrak{C}_{\lambda}$ and Lusztig’s asymptotic algebra
Fix bijections $\phi_{ij}:\Gamma_{i}\cap\Gamma_{j}^{-1}\to\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$, as in [34, §2.3]. Write $E_{ij}(a)$ for the matrix with $a$ in the $(i,j)$-position and 00 elsewhere. The following result of Xi reduces the understanding of $\mathcal{J}_{\lambda}$ to understanding the ring $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$.
Theorem 7.8 ([34, Theorems 2.3.2 and 8.4.2]).
@@ 2. $(b)$
By Theorem 7.2 we have $\mathfrak{c}_{\lambda}(d_{i})=E_{ii}$. Let $w_{ij}=\phi^{-1}_{ij}(\mathsf{w}_{\lambda^{\prime}})\in\Gamma_{i}\cap\Gamma_{j}^{-1}$. By [34, Lemma 2.3.1] we have $w_{ii}=d_{i}$ and $w_{ji}=w_{ij}^{-1}$. Since $E_{ij}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})E_{ji}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})=E_{ii}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}}^{2})=E_{ii}(\mathsf{t}_{\mathsf{w}_{\lambda^{\prime}}})$ it follows (under the isomorphism in Theorem 7.8(2)) that $\mathsf{t}_{w_{ij}}\mathsf{t}_{w_{ij}^{-1}}=\mathsf{t}_{d_{i}}.$ By Theorem 7.2 we have $\mathfrak{c}_{\lambda}(w_{ij})=a(\zeta_{\lambda})E_{ij}$ for some $a(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]$. Then the above equation, and the isomorphism from Theorem 5.11, gives (using (4.5)) $E_{ii}=\mathfrak{c}_{\lambda}(d_{i})=\mathfrak{c}_{\lambda}(w_{ij})\mathfrak{c}_{\lambda}(w_{ij}^{-1})=a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})E_{ii},$ and hence $a(\zeta_{\lambda})a(\zeta_{\lambda}^{-1})=1$. Writing $a(\zeta_{\lambda})=\sum_{\gamma}c_{\gamma}\zeta_{\lambda}^{\gamma}$ (with $c_{\gamma}\in\mathbb{Z}$ and the sum over $\gamma\in P^{(\lambda)}$) it follows that $\sum_{\gamma}c_{\gamma}^{2}=1$, and hence $a(\zeta_{\lambda})=\epsilon_{ij}\zeta_{\lambda}^{\gamma_{ij}}$ for some $\gamma_{ij}\in P^{(\lambda)}$ and $\epsilon_{ij}=\pm 1$. Thus the elements $w_{ij}$ satisfy $\mathfrak{c}_{\lambda}(w_{ij})=\epsilon_{ij}\zeta_{\lambda}^{\gamma_{ij}}E_{ij},$ with $\epsilon_{ji}=\epsilon_{ij}$ and $\gamma_{ji}=-\gamma_{ij}$ (with $\epsilon_{ii}=1$).
@@
and so $\mathfrak{c}_{\lambda}(\mathsf{w}_{1i})A\mathfrak{c}_{\lambda}(\mathsf{w}_{j1})=a(\zeta_{\lambda})E_{11}$, and hence $a(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$. So $\mathcal{R}_{ij}\subseteq\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$. On the other hand, $D^{-1}\mathfrak{c}_{\lambda}(w_{i1})\mathfrak{c}_{\lambda}(\mathsf{m}_{\gamma})\mathfrak{c}_{\lambda}(w_{1j})D=\mathfrak{s}_{\gamma}(\zeta_{\lambda})E_{ij},$ and so $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}\subseteq\mathcal{R}_{ij}$, completing the proof of the claim. Thus $D^{-1}\mathfrak{C}_{\lambda}D=\mathrm{Mat}_{N_{\lambda}}(\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}})$.
@@
and so $\mathfrak{c}_{\lambda}(\mathsf{w}_{1i})\mathfrak{c}_{\lambda}(w)\mathfrak{c}_{\lambda}(\mathsf{w}_{j1})=\epsilon_{w}\mathfrak{c}_{\lambda}(\mathsf{m}_{h_{\lambda}(w)})$. Thus $\mathsf{t}_{\mathsf{w}_{1i}}\mathsf{t}_{w}\mathsf{t}_{\mathsf{w}_{j1}}=\epsilon_{w}\mathsf{t}_{\mathsf{m}_{h_{\lambda}(w)}}.$ It is known that the structure constants of $\mathcal{J}_{\lambda}$ are nonnegative (see [19, (3.1.1)]), and hence $\epsilon_{w}=1$. Thus we have $D^{-1}\mathfrak{c}_{\lambda}(w)D=\mathfrak{s}_{h_{\lambda}(w)}(\zeta_{\lambda})$ for all $w\in\Delta_{\lambda}$. ∎
@@
Theorem 7.9 gives rise to a balanced system of cell representations, in the sense of [15, Definition 1.5].
@@
We verify properties $\mathbf{B1}$$\mathbf{B6}$ from [15, Definition 1.5]. Property $\mathbf{B1}$ is the killing property (Theorem 4.3), and property $\mathbf{B2}$ is boundedness (Theorem 4.7, with bounds $\mathbf{a}_{\lambda}=\ell(\mathsf{w}_{\lambda^{\prime}})$). Property $\mathbf{B6}$ (monotonicity of the bounds) is then immediate from Lemma 1.3. Property $\mathbf{B3}$ is the statement that $\pi_{\lambda}$ recognises precisely $\Delta_{\lambda}$ (Theorem 5.10), and property $\mathbf{B4}$ (freeness of the leading matrices over $\mathbb{Z}$) follows from the freeness of the Schur functions and Theorem 7.9. Finally, if $z\in\Delta_{\lambda}$ then $z\in\Gamma_{i}\cap\Gamma_{j}^{-1}$ for some $1\leq i,j\leq N_{\lambda}$ and by Theorem 7.9 we have $\mathfrak{c}_{\lambda}(z)\mathfrak{c}_{\lambda}(d_{j})=\mathfrak{c}_{\lambda}(z)$, verifying $\mathbf{B5}$. ∎
@@
Conjugation by $D$ in the above theorem amounts to choosing a basis associated to a (signed) “fundamental domain” for the action of $\mathsf{T}_{\lambda}$ on $\mathcal{A}_{\lambda}$ different from the standard fundamental domain ${{}^{\lambda}}W$. See [13, Section 4.5] for details. Specifically, the basis chosen is $\{\xi_{\lambda}\otimes\epsilon_{i}X_{\tau_{\gamma_{i}}^{-1}u_{i}}\mid 1\leq i\leq N_{\lambda}\}$ where ${{}^{\lambda}}W=\{u_{i}\mid 1\leq i\leq N_{\lambda}\}$ is ordered as in Theorem 7.2.
@@
$w_{12}$ $=s_{1}s_{3}s_{2}$ $w_{13}$ $=s_{1}s_{3}s_{0}\sigma$ $w_{14}$ $=s_{1}s_{3}s_{2}\sigma$ $w_{15}$ $=s_{1}s_{3}\sigma$ $w_{16}$ $=s_{1}s_{3}s_{2}\sigma^{2}$
$w_{23}$ $=s_{2}s_{1}s_{3}s_{0}\sigma$ $w_{24}$ $=s_{2}s_{1}s_{3}s_{2}\sigma$ $w_{25}$ $=s_{2}s_{1}s_{3}\sigma$ $w_{26}$ $=s_{2}s_{1}s_{3}s_{2}\sigma^{2}$ $w_{34}$ $=s_{3}s_{2}s_{0}s_{1}$
$w_{35}$ $=s_{3}s_{2}s_{0}$ $w_{36}$ $=s_{3}s_{2}s_{0}s_{1}\sigma$ $w_{45}$ $=s_{1}s_{2}s_{0}$ $w_{46}$ $=s_{1}s_{2}s_{0}s_{1}\sigma$ $w_{56}$ $=s_{2}s_{0}s_{1}\sigma$
@@ 2. $(2)$
It is sufficient to prove (1), because (2) follows by applying (1) to $(t_{\gamma}u)^{-1}=t_{-u^{-1}\gamma}u^{-1}$. By counting hyperplanes separating $\mathcal{A}_{0}$ from $t_{\gamma}u\mathcal{A}_{0}$, we have (see [25, (2.4.1)])
@@ A.4 Monotonicity of $\ell(\mathsf{m}_{\gamma})$ with respect to $\preccurlyeq_{\lambda}$
and similarly for $L(\gamma^{\prime})$. Write $\gamma^{\prime}+Q_{\lambda}=\gamma+\epsilon+Q_{\lambda}$ with $\epsilon\in Q^{\lambda}_{+}$. From Lemma 2.16 we have $\epsilon=\sum_{i=1}^{r(\lambda)}d_{i}\tilde{e}_{i}$ where $\sum_{i\in B_{p}}d_{i}=0$ for all $1\leq p\leq r$ and $d_{1}+\cdots+d_{i}\geq 0$ for all $1\leq i\leq r(\lambda)$. Thus $\epsilon=\epsilon_{1}+\cdots+\epsilon_{r}$ where $\epsilon_{p}=\sum_{i\in B_{p}}d_{i}\tilde{e}_{i}\in Q^{\lambda}_{+}$, and for any $1\leq p\leq r$ we have $\gamma+\epsilon_{p}+Q_{\lambda}\in(P/Q_{\lambda})_{+}$ with $\gamma+Q_{\lambda}\preccurlyeq_{\lambda}\gamma+\epsilon_{p}+Q_{\lambda}\preccurlyeq_{\lambda}\gamma^{\prime}+Q_{\lambda}$. Thus it suffices to prove the result in the case $\gamma^{\prime}+Q_{\lambda}=\gamma+\epsilon_{p}+Q_{\lambda}$ for some $1\leq p\leq r$ (that is, $\gamma+Q_{\lambda}$ and $\gamma^{\prime}+Q_{\lambda}$ only differ in one block), and then by well known properties of the dominance order on the symmetric group (see, for example, [32, Corollary 2.7]) it suffices to prove the result when $\epsilon_{p}=\tilde{e}_{i}-\tilde{e}_{j}$ for some $i,j\in B_{p}$ with $i<j$.